# A Construction of non-degenerate $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{n}$

### v1

@@ Abstract.
We discover an explicit construction of non-degenerate $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{n},n\geq 3$, using a variant of ellipsoidal coordinates on $\mathbb{R}^{n}$. The branching set of these examples is a codimension-$2$ ellipsoid. Moreover, the graph of the related $\mathbb{Z}_{2}$-harmonic one form in $T^{*}\mathbb{R}^{n}$, can be obtained as a certain limit of a specific sequence of Lawlor’s necks in $\mathbb{C}^{n}=T^{*}\mathbb{R}^{n}$.
@@ 1. Introduction
Let $(M^{n},g),n\geq 3$ be an oriented Riemannian manifold and $\Sigma\subset M$ be an oriented codimension-$2$ submanifold. Let $L$ be a flat $\mathbb{R}$ bundle over $M\setminus\Sigma$, such that any small loop linking along $\Sigma$ has monodromy $-1$ and $(E,\nabla)$ be a vector bundle equipped with a connection. A section $\alpha\in\Gamma(M\setminus\Sigma,E\otimes L)$ is called a $\mathbb{Z}_{2}$-harmonic section if $\alpha,\nabla\alpha\in L_{loc}^{2}$ and satisfies the harmonic equation
$\nabla^{*}\nabla\alpha=0,\text{ on }M\setminus\Sigma.$
Here, the derivatives are taken with respect to the flat structure on $L$. Of particular interest are the cases where $E$ is a trivial line bundle with trivial connection, cotangent bundle with Levi-Civita connection, or spinor bundle with spin connection when $M$ is spin. In these cases, the harmonic sections are referred to as $\mathbb{Z}_{2}$-harmonic functions, one-forms and spinors, respectively. Alternatively, $\mathbb{Z}_{2}$-harmonic sections can be viewed as a section on a double branched cover, which is anti-invariant under the covering involution.
Those multi-valued harmonic sections play important roles in the compactification problem in gauge theory [9], and in the field of special holonomy and calibrated geometry [2, 5].
In this paper, we focus mostly on $\mathbb{Z}_{2}$-harmonic functions. Let $N$ be the normal bundle of $\Sigma$, and use the inverse of the exponential to define the map $\zeta$ from a tubular neighborhood of $\Sigma$ to $N$ in $(M,g)$. Use co-orientation we can regard $N$ as a complex line bundle over $\Sigma$. If $\sigma$ is a section of the dual normal bundle $N^{-1}$, we get a complex-valued function $\langle\sigma,\zeta\rangle$, which we will simply denote as $\sigma\zeta$. The monodromy around $\Sigma$ defines a square root $N^{1/2}$ with dual $N^{-1/2}$. Similarly, we can define a complex-valued function
@@
for any half-integer $p$. According to Donaldson [1], any $\mathbb{Z}_{2}$-harmonic function $f$ admits an asymptotic behavior near $\Sigma$
@@
Donaldson showed that if $A\equiv 0$ and $B$ are nowhere-vanishing, then $f$ admits a good deformation theory. Such $\mathbb{Z}_{2}$-harmonic functions are called non-degenerate, and will appear in gluing construction, for example [6, 12].
Here, we will only consider harmonic functions on $\mathbb{R}^{n},n\geq 3$, with branching set $\Sigma$ being a compact smooth embedded submanifold. Outside a large ball that contains $\Sigma$, $L$ is a trivial line bundle. In other words, we can pick a single-valued branch of the $\mathbb{Z}_{2}$-harmonic function. Moreover, different choices differ by a sign. In this setting we can regard a $\mathbb{Z}_{2}$-harmonic function $f$ as a harmonic function outside a large ball. Suppose $f$ has polynomial growth, then classical theory implies:
@@
where $P$ is a harmonic polynomial. Weifeng Sun [7] and later Andriy et al [8]. showed that when $\Sigma$ is fixed, there is a $1$-$1$ correspondence between $\mathbb{Z}_{2}$ harmonic functions and harmonic polynomials $P$ related by Equation 1.1. In their existence result, little is known about the asymptotic behavior. On the other hand, recently Donaldson found a family of $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{3}$ with branching set being codimension-$2$ ellipses using twistor theory [3]. The goal of this paper is to use a variant of ellipsoidal coordinates to give an explicit construction of non-degenerate $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{n}$, which recovers Donaldson’s construction when $n=3$.
@@
Here, let $S(y)=\prod_{i=1}^{n-1}(y+h_{i}^{2})$, $a_{i}$ are constants given by
@@
Let $P$ be a degree-two homogeneous harmonic polynomial with the origin being its only critical point, whose index is $n-1$. Let $C_{n}$ be a positive constant. Then there exists a non-degenerate $\mathbb{Z}_{2}$-harmonic function $f$ constructed in Theorem 1.1, such that we can pick a single-valued branch of $f$, on which
@@
Furthermore, when $C_{n}$ goes to zero, the ellipsoid scales down to 00 and $f$ converges to $P$ outside any ball that contains the ellipsoid.
@@
The corresponding $\mathbb{Z}_{2}$-harmonic one-form $\mathrm{d}f$ can be viewed as a higher-dimensional analogue of $\mathrm{Re}(\sqrt{z^{2}-\epsilon^{2}}\mathrm{d}z)$ on $\mathbb{C}$.
In a slightly different vein, in the compact setting, a $\mathbb{Z}_{2}$-harmonic one form is the infinitesimal branched deformation of special Lagrangians [5]. In a suitable setting, if we regard $T^{*}\mathbb{R}^{n}$ as $\mathbb{C}^{n}$, Lawlor’s necks can be viewed as a $2$-valued graph. More precisely, there exists an $\mathbb{Z}_{2}$-potential function $F$ on $\mathbb{R}^{n}$, such that Lawlor’s necks is the $2$-valued graph $\mathrm{d}F$ from $\mathbb{R}^{n}$ to $T^{*}\mathbb{R}^{n}$. A version of the infinitesimal branched deformation in our case is the following.
@@ Acknowledgement
The author is grateful to his Ph.D. advisor Claude LeBrun and Simon Donaldson for their suggestions, discussion, and encouragement. He would also like to thank Jingrui Cheng, Dylan Galt, Siqi He, and Yusen Xia for fruitful discussion and helpful comments. This work is partially supported by Simons Foundation.
@@ 2. Ellipsoidal coordinates on $\mathbb{R}^{n}$
Our construction is motivated by ellipsoidal harmonics in classical electromagnetism. These functions are solutions to the Laplacian equation on $\mathbb{R}^{n}$ under ellipsoidal coordinates, a coordinate system that is orthogonal and is well-defined on an open dense subset in $\mathbb{R}^{n}$. The key to our construction is the discovery that a variant of the ellipsoidal coordinates can be extended to orthogonal coordinates on a double cover of $\mathbb{R}^{n}$ with the branching set being an ellipsoid.
In this section, we will give a brief review on the ellipsoidal coordinates and its variant in $\mathbb{R}^{n},n\geq 3$. Our main reference for this section is [4]. Consider a family of hypersurfaces $C_{i},1\leq i\leq n$ defined by
@@
The parameters $(\lambda_{i})$ define an orthogonal coordinates system in $\mathbb{R}^{n}$, which is known as ellipsoidal coordinate. In this coordinate, the Euclidean metric is
@@
Let $\sigma_{i,k}(\boldsymbol{y}),1\leq i\leq n-1$ be the $k$-th symmetric polynomial of $n-2$ variables
@@ 2.1. Branched cover and modified ellipsoidal coordinate
Here, $h_{i+1}^{2}<\mu_{i}^{2}<h_{i}^{2},i<n-1$, and $00<\mu_{n-1}^{2}<h_{n-1}^{2},\mu_{n}^{2}>$ 0. After the change of parameters, Equation 2.2 becomes:
@@
If we view $(\theta_{1},\cdots,\theta_{n-2})$ as spherical coordinates $\Theta$, that parametrize an ellipsoid, then the Equation LABEL:S_2_Eqn_Coor_3 defines a smooth map
@@
We call $\boldsymbol{x}(\Theta,\mu_{n-1},\mu_{n})$ as modified ellipsoidal coordinate, here the parameters $h_{i}$ are allowed to satisfy $0<h_{i+1}\leq h_{i}$.
It is easy to see that $\boldsymbol{x}$ is generically a $2$-$1$ map, such that
@@
Moreover, we will have the following proposition
@@
The coordinates map $\boldsymbol{x}$ defines an orthogonal coordinate system on the double cover over $\mathbb{R}^{n}$, with branching set
@@
To begin, we first consider the case when $h_{i+1}<h_{i}$ and then derive a formula that can extend to $h_{i+1}\leq h_{i}$. We define $\sigma_{k}(\boldsymbol{h}),\sigma_{j,k}(\boldsymbol{h}),\sigma_{k}(\boldsymbol{\mu})$ as the symmetric polynomials with respect to the multiples
@@
The symmetric polynomials $\sigma_{k}(\boldsymbol{\mu})$ are linear combinations of constant and $x_{i}^{2}$. In particular, we have
@@
Apply Lemma 2.1, we prove the lemma. ∎
@@ 3. Construction
Our strategy to the construction is that we will find solutions to the Laplacian equation, which takes the form
@@ 3.1. Separation of Variables
We will use Equation 2.5 to write down the Laplacian operator under $\mu_{i}$’s coordinate. Let $S(y)=\prod_{j=1}^{n-1}(h_{j}^{2}+y)$, then the Laplacian operator $\Delta$ is
@@
Let $Q(y)=\sum_{k=0}^{n-2}Q_{k}y^{k}$ be any polynomial with degree less than $n-1$ and $y_{j}=\mu_{j}^{2},y_{n}=-\mu_{n}^{2}$. It follows from Lemma 2.1 that
@@
We define further that
@@
The equation $\mathscr{L}_{j,Q}f=0$ are identically the same as
@@
We will begin with finding quadratic polynomial solutions to Equation 3.1, which is closely related to harmonic polynomials with quadratic growth in $\mathbb{R}^{n}$. Assume $f=\mu^{2}-p$, then $f$ is a solution is equivalent to
$02S(-\mu^{2})-2\mu^{2}S^{\prime}(-\mu^{2})+Q(\mu^{2})(\mu^{2}-p)=$ 0 .
@@
then we can solve $Q_{i},i=0,\cdots,n-2$ inductively. The equation $S(-y)=0$ has $n-1$ distinct positive roots, and henceforth $S(-y)-yS^{\prime}(-y)=0$ has $n-1$ distinct positive solutions, which we denote as $p_{i}$. In particular, we can arrange $p_{i}$ such that
@@
is a solution to the Laplacian equation, which is known as an ellipsoidal harmonic of degree $2$ [4]. The Lemma 2.6 implies that each $P_{i}$ is a harmonic polynomial of the forms
@@ 3.2. Construction of non-degenerate $\mathbb{Z}_{2}$-harmonic function
Follow from Proposition 2.5 if $f=f(\Theta,\mu_{n-1},\mu_{n})$ is an odd function with respect to $(\mu_{n},\mu_{n-1})$, then $f$ defines a multivalued function on $\mathbb{R}^{n}$ with monodromy $-1$ around the branching set $E_{\boldsymbol{h}}$. Recall that $S(y)=\prod_{j=1}^{n-1}(y+h_{j}^{2})$, and we define
@@
Those harmonic functions are odd functions with respect to $(\mu_{n},\mu_{n-1})$, and henceforth are $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{n}$. Moreover, $|f_{0}|$ are bounded and $|f_{2,i}|$ grow no faster than quadratic in $\mathbb{R}^{n}$.
@@
is strictly less than 00 on the branching set $E_{\boldsymbol{h}}$.

### v2

@@ Abstract.
We discover an explicit construction of non-degenerate $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{n},n\geq 3$, using a variant of ellipsoidal coordinates on $\mathbb{R}^{n}$. The branching set of these examples is a codimension-$2$ ellipsoid, providing the first known family of non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms on $\mathbb{R}^{n}$ with compact branching sets. Moreover, the graph of the related $\mathbb{Z}_{2}$-harmonic one form in $T^{*}\mathbb{R}^{n}$ can be obtained as a certain limit of a specific sequence of Lawlor’s necks in $\mathbb{C}^{n}=T^{*}\mathbb{R}^{n}$.
@@ 1. Introduction
Let $(M^{n},g),n\geq 2$ be an oriented Riemannian manifold, and $\Sigma\subset M$ be an oriented codimension-$2$ submanifold. Let $L$ be a flat $\mathbb{R}$ bundle over $M\setminus\Sigma$, such that any small loop linking along $\Sigma$ has monodromy 1-1 and $(E,\nabla)$ be a vector bundle equipped with a connection. A section $\alpha\in\Gamma(M\setminus\Sigma,E\otimes L)$ is called a $\mathbb{Z}_{2}$-harmonic section if $\alpha,\nabla\alpha\in L_{loc}^{2}$ and satisfies the harmonic equation
$\nabla^{*}\nabla\alpha=0,\text{ on }M\setminus\Sigma,$
where the derivatives are taken with respect to the flat structure on $L$. Of particular interest are the cases where $E$ is a trivial line bundle with a trivial connection, a cotangent bundle with a Levi-Civita connection, or a spinor bundle with a spin connection when $M$ is spin. In these cases, the harmonic sections are referred to as $\mathbb{Z}_{2}$-harmonic functions, $1$-forms, and spinors, respectively. Alternatively, $\mathbb{Z}_{2}$-harmonic sections can be viewed as a section on a double branched cover, which is anti-invariant under the covering involution.
Multi-valued harmonic sections play important roles in the compactification problems in gauge theory [14, 12, 10, 6] and in the field of special holonomy and calibrated geometry [3, 7]. Despite its importance, a general existence theorem for $\mathbb{Z}_{2}$-harmonic sections is impossible, as it involves the choice of $\Sigma$, which is intrinsically a nonlinear problem. Research in this direction has been largely example-driven. Several examples, both compact and non-compact, have been studied previously by many people [1, 8, 9, 13, 15, 16].
In this paper, we provide an explicit construction for non-degenerate $\mathbb{Z}_{2}$-harmonic functions when $M=\mathbb{R}^{n},n\geq 3$ and $\Sigma$ are codimension-$2$ ellipsoids. Our motivation for the construction originates from Donaldson’s program [2] on the adiabatic limit of $G_{2}$ manifolds with coassociative $K3$ fibrations and branched maximal sections, which are a nonlinear analog of non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms, appearing as the potential adiabatic limiting description for the $G_{2}$ structure. The examples we construct serve as geometric models to resolve the problem when the branched maximal sections do not avoid the (2)(-2)-classes. We leave the further discussion in this direction to Section 1 in [19].
To begin, we focus on $\mathbb{Z}_{2}$-harmonic functions. Let $N$ be the normal bundle of $\Sigma$, and use the inverse of the exponential to define the map $\zeta$ from a tubular neighborhood of $\Sigma$ to $N$ in $(M,g)$. Using co-orientation, we can regard $N$ as a complex line bundle over $\Sigma$. If $\sigma$ is a section of the dual normal bundle $N^{-1}$, we obtain a complex-valued function $\langle\sigma,\zeta\rangle$, which we will simply denote as $\sigma\zeta$. The monodromy around $\Sigma$ defines a square root $N^{1/2}$ with dual $N^{-1/2}$. Similarly, we can define a complex-valued function
@@
for any half-integer $p$. According to Donaldson [2], any $\mathbb{Z}_{2}$-harmonic function $f$ admits asymptotic behavior near $\Sigma$
@@
Donaldson showed that if $A\equiv 0$ and $B$ are nowhere-vanishing, then $f$ admits a good deformation theory. Such $\mathbb{Z}_{2}$-harmonic functions are called non-degenerate and will appear in gluing construction, for example, [8, 19].
Here, we will only consider harmonic functions on $\mathbb{R}^{n},n\geq 3$, with the branching set $\Sigma$ being a compact smooth embedded submanifold. Outside a large ball that contains $\Sigma$, $L$ is a trivial line bundle. In other words, we can pick a single-valued branch of the $\mathbb{Z}_{2}$-harmonic function. Moreover, different choices differ by a sign. In this setting, we can regard a $\mathbb{Z}_{2}$-harmonic function $f$ as a harmonic function outside a large ball. Suppose $f$ has polynomial growth, then classical theory implies:
@@
where $P$ is a harmonic polynomial. Weifeng Sun [11] and later Haydys, Mazzeo, Takahashi [13] showed that when $\Sigma$ is fixed, there is a 11-11 correspondence between $\mathbb{Z}_{2}$ harmonic functions and harmonic polynomials $P$ related by Equation 1.1. In their existence result, little is known about the asymptotic behavior. On the other hand, recently, Donaldson found a family of $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{3}$ with the branching set being codimension-$2$ ellipses using twistor theory [4]. The goal of this paper is to use a variant of ellipsoidal coordinates to give an explicit construction of non-degenerate $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{n}$, which recovers Donaldson’s construction when $n=3$.
@@
Here, let $S(y)=\prod_{i=1}^{n-1}(y+h_{i}^{2})$ and $a_{i}$ be constants given by
@@
Let $P$ be a degree-two homogeneous harmonic polynomial with the origin as its only critical point, whose index is n1n-1. Let $C_{n}$ be a positive constant. Then there exists a unique non-degenerate $\mathbb{Z}_{2}$-harmonic function $f$ constructed in Theorem 1.1, such that we can pick a single-valued branch of $f$, on which
@@
Furthermore, when $C_{n}$ goes to zero, the ellipsoid scales down to 0, and $f$ converges to $P$ outside any ball that contains the ellipsoid.
@@
The corresponding $\mathbb{Z}_{2}$-harmonic $1$-form $\mathrm{d}f$ can be viewed as a higher-dimensional analog of $\mathrm{Re}(\sqrt{z^{2}-\epsilon^{2}}\mathrm{d}z)$ on $\mathbb{C}$. Roughly speaking, this family of $1$-forms serves as a local model of a family of non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms with a shrinking branching set. In [19], we develop a gluing theorem that glues these examples to the regular zeros of harmonic $1$-forms on compact manifolds, thereby producing many new examples of $\mathbb{Z}_{2}$-harmonic $1$-forms on compact manifolds.
In a related direction, in the compact setting, a $\mathbb{Z}_{2}$-harmonic one form is the infinitesimal branched deformation of special Lagrangians [7]. In a suitable setting, if we regard $T^{*}\mathbb{R}^{n}$ as $\mathbb{C}^{n}$, Lawlor’s necks can be viewed as a $2$-valued graph. More precisely, there exists an $\mathbb{Z}_{2}$-potential function $F$ on $\mathbb{R}^{n}$, such that Lawlor’s necks is the $2$-valued graph $\mathrm{d}F$ from $\mathbb{R}^{n}$ to $T^{*}\mathbb{R}^{n}$. A version of the infinitesimal branched deformation in our case is the following.
@@ Acknowledgement
The author is grateful to his Ph.D advisor, Prof. LeBrun, as well as to Prof. Donaldson for their valuable discussions and encouragement. He would also like to thank Jingrui Cheng, Dylan Galt, Siqi He, and Yusen Xia for fruitful discussions and helpful comments. This work is partially supported by Simons Foundation.
@@ 2. Ellipsoidal coordinates on $\mathbb{R}^{n}$
Our construction is motivated by ellipsoidal harmonics in classical electromagnetism. These functions are solutions to the Laplacian equation on $\mathbb{R}^{n}$ under ellipsoidal coordinates, a coordinate system that is orthogonal and well-defined on an open dense subset in $\mathbb{R}^{n}$. The key to our construction is the discovery that a variant of the ellipsoidal coordinates can be extended to orthogonal coordinates on a double cover of $\mathbb{R}^{n}$ with the branching set being an ellipsoid.
In this section, we will give a brief review of the ellipsoidal coordinates and their variant in $\mathbb{R}^{n},n\geq 3$. Our main reference for this section is [5]. Consider a family of hypersurfaces $C_{i},1\leq i\leq n$ defined by
@@
The parameters $(\lambda_{i})$ define an orthogonal coordinate system in $\mathbb{R}^{n}$, which is known as ellipsoidal coordinates. In this coordinate, the Euclidean metric is
@@
Let $\sigma_{i,k}(\boldsymbol{y}),1\leq i\leq n-1$ be the $k$-th symmetric polynomial of n2n-2 variables
@@ 2.1. Branched cover and modified ellipsoidal coordinate
Here, $h_{i+1}^{2}<\mu_{i}^{2}<h_{i}^{2},i<n-1$, and $0<\mu_{n-1}^{2}<h_{n-1}^{2},\mu_{n}^{2}>0$. After the change of parameters, Equation 2.2 becomes:
@@
If we view $(\theta_{1},\cdots,\theta_{n-2})$ as spherical coordinates $\Theta$ that parametrize an ellipsoid, then the Equation LABEL:S_2_Eqn_Coor_3 defines a smooth map
@@
We call $\boldsymbol{x}(\Theta,\mu_{n-1},\mu_{n})$ a modified ellipsoidal coordinate, here the parameters $h_{i}$ are allowed to satisfy $0<h_{i+1}\leq h_{i}$.
It is easy to see that $\boldsymbol{x}$ is generically a 22-11 map, such that
@@
Moreover, we will have the following proposition:
@@
The coordinates map $\boldsymbol{x}$ defines an orthogonal coordinate system on the double cover over $\mathbb{R}^{n}$, with a branching set
@@
To begin, we first consider the case when $h_{i+1}<h_{i}$ and then derive a formula that can be extended to $h_{i+1}\leq h_{i}$. We define $\sigma_{k}(\boldsymbol{h}),\sigma_{j,k}(\boldsymbol{h}),\sigma_{k}(\boldsymbol{\mu})$ as the symmetric polynomials with respect to the multiples
@@
The symmetric polynomials $\sigma_{k}(\boldsymbol{\mu})$ are linear combinations of constants and $x_{i}^{2}$. In particular, we have
@@
Applying Lemma 2.1, we prove the lemma. ∎
@@ 3. Construction
Our strategy for the construction is to find solutions to the Laplacian equation, which takes the form
@@ 3.1. Separation of Variables
We will use Equation 2.5 to write down the Laplacian operator under $\mu_{i}$’s coordinates. Let $S(y)=\prod_{j=1}^{n-1}(h_{j}^{2}+y)$, then the Laplacian operator $\Delta$ is
@@
Let $Q(y)=\sum_{k=0}^{n-2}Q_{k}y^{k}$ be any polynomial with degree less than n1n-1 and $y_{j}=\mu_{j}^{2},y_{n}=-\mu_{n}^{2}$. It follows from Lemma 2.1 that
@@
We further define that
@@
The equation $\mathscr{L}_{j,Q}f=0$ is identically the same as
@@
We will begin by finding quadratic polynomial solutions to Equation 3.1, which is closely related to harmonic polynomials with quadratic growth in $\mathbb{R}^{n}$. Assume $f=\mu^{2}-p$, then $f$ is a solution, which is equivalent to
$2S(-\mu^{2})-2\mu^{2}S^{\prime}(-\mu^{2})+Q(\mu^{2})(\mu^{2}-p)=0.$
@@
then we can solve $Q_{i},i=0,\cdots,n-2$ inductively. The equation S(y)=0S(-y)=0 has n1n-1 distinct positive roots, and henceforth $S(-y)-yS^{\prime}(-y)=0$ has n1n-1 distinct positive solutions, which we denote as $p_{i}$. In particular, we can arrange $p_{i}$ such that
@@
is a solution to the Laplacian equation, which is known as an ellipsoidal harmonic of degree $2$ [5]. The Lemma 2.6 implies that each $P_{i}$ is a harmonic polynomial of the form
@@ 3.2. Construction of non-degenerate $\mathbb{Z}_{2}$-harmonic function
Follow from Proposition 2.5 if $f=f(\Theta,\mu_{n-1},\mu_{n})$ is an odd function with respect to $(\mu_{n},\mu_{n-1})$, then $f$ defines a multivalued function on $\mathbb{R}^{n}$ with monodromy 1-1 around the branching set $E_{\boldsymbol{h}}$. Recall that $S(y)=\prod_{j=1}^{n-1}(y+h_{j}^{2})$, and we define
@@
Those harmonic functions are odd functions with respect to $(\mu_{n},\mu_{n-1})$, and hence, they are $\mathbb{Z}_{2}$-harmonic functions on $\mathbb{R}^{n}$. Moreover, $|f_{0}|$ is bounded, and $|f_{2,i}|$ grows no faster than quadratic in $\mathbb{R}^{n}$.
@@
is strictly less than 0 on the branching set $E_{\boldsymbol{h}}$.